19a-3(n+3)=9 whats n??

Answer by tutorcecilia(2152)

19a-3(n+3)=9 solve for n
19a-3n-9=9
19a-3n-9+9=9+9
19a-3n-19a=18-19a
-3n=18-19a
-3n/-3=18-19a/-3
n=(-18+9a)/3